A o/w emulsion contains mineral oil (density 0.9 kg/L) and aqueous medium (density 1.05 kg/L, viscocity 0.5 poise. If the droplets have an average diameter of 5 μm, what is the velocity (cm/day) of creaming in 24 hours, according to stokes' equation?
Answer #1

The creaming ratio for o/w emulsions can be estimated from the Stokes equation:

2r^(p-polg -- U = 9n

where, υ is the creaming (settling) rate, r is the droplet radius, ρ is the density of the droplet, ρo is the density of the dispersion medium, η is the viscosity of the dispersion medium (continuous phase) and g is the local acceleration due to gravity.

If we put our values in the equation, we get (I've converted Poise to Pa.s for easier unit conversion later, 10 poise = 1 Pa.s):

U =- 2. (2.5x10-6m)?(0.9kg – 1.054) - 9.81 32 = 4.087. 10-11 9.0,05Pa.

This can be converted to cm/day as:

um 5 100cm 1m 3600s 1h 24h 1day = 3.53.10-4 day